OpenStudy (anonymous):

What is the difference between rational/irrational number, integer, real number, complex number and natural number ?

OpenStudy (anonymous):

rational- number that can be expressed as fraction. irrational- PI cant be expressed as fraction. Integer- Any whole real number Real number- any normal number-- PI, 1, 1.25 etc complex- a number with an imaginary component- 1+i natural number- nonnegative integer

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

PI is prime integer ? what is a imaginary component ?

OpenStudy (anonymous):

PI is not an integer. Imaginary component just has the letter i in it like 1+2i

OpenStudy (anonymous):

http://www.youtube.com/watch?v=57unv_eJC14

OpenStudy (anonymous):

in other words an "imaginary number" i=sqrt(-1) complex number= a+bi where a and b are any real number

OpenStudy (anonymous):

watch them all u will understand more lol @rox13kh

OpenStudy (anonymous):

thx for the link @Yahoo!

OpenStudy (anonymous):

welcome>>>

OpenStudy (anonymous):

getting the hang of it i guess @ivan514 thx

OpenStudy (anonymous):

so PI is a prime number ?

OpenStudy (anonymous):

mmm can others give answers too ? just to make sure

OpenStudy (ash2326):

Hey rox13kh, did you understand this?

OpenStudy (anonymous):

I still don't get rational and complex

OpenStudy (ash2326):

A rational number is of the form \[\frac{a}{b}\] where b can't be 0 a can be any whole number like \(\frac{2}{4}\) or \(\frac{2}{1}\)

OpenStudy (anonymous):

thx so irrational are those with denominators 0 and decimal ? ? how about complex

OpenStudy (ash2326):

nope, irrational are numbers which can't be expressed as a/b like \(\sqrt 2\) or \(\pi\)

OpenStudy (anonymous):

got it. what are complex numbers ?

OpenStudy (anonymous):

a wait.. can't pi be expressed as 22/7 ? ?

OpenStudy (ash2326):

you know if we have \[\sqrt{-1}\] it's represented as i so \[i=\sqrt{-1}\] this is called an imaginary number A number of the form \[a+ib\] is a complex no. Do you get till here?

OpenStudy (anonymous):

yeah2 u can continue

OpenStudy (ash2326):

here a is the real part and b is the imaginary part You know it's amazing that every number is a complex number, 2 is a complex no. with 0 imaginary part \(\sqrt {-1} \) is a complex no. with 0 real part

OpenStudy (anonymous):

got it thx a lot ^^

OpenStudy (ash2326):

You;re welcome. @rox13kh anytime you need help here, do let me know

OpenStudy (anonymous):

sure